# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ A.4 $P$-Targeted Kernel Stein Discrepancy
This section states preliminary results which will be needed to prove Theorem 4. The main tool we use is the following generalized definition of KSD, which was originally proposed in Shi and Mackey, (2024, Definition 4).
@@
  • Lemma 15 shows that the standard KSD can be written as a $P$-KSD.
  • Lemma 16 shows that $\mathbb{S}_{P}(Q,R)$ is equivalent to an MMD, and that it admits a closed form involving expectations over $Q$ and $R$.
  • Lemma 17 shows that the $P$-KSD-projection of a probability measure onto a (standard) KSD-ball centered at $P$ admits a closed-form expression.
  • Lemma 18 restates Tolstikhin et al., (2017, Proposition A.1), which is a McDiarmid-type inequality for MMD. It also applies to $P$-KSD, as an $P$-KSD is also an MMD. We will use it in Appendix D to derive a robust-KSD test that is well-calibrated for finite samples. Other deviation bounds instead of the McDiarmid bound could also be used to construct a similar test;see Remark 24.1 for a discussion.
The MMD has already been studied extensively in the context of both Bayesian and frequentist robust parameter estimation; see, e.g., Briol et al., (2019); Chérief-Abdellatif and Alquier, (2020, 2022); Dellaporta et al., (2022); Dellaporta and Damoulas, (2023); Alquier and Gerber, (2024). A key assumption to ensure robustness in all of these papers is that the kernel is assumed to be bounded, and this is exactly what we are able to achieve with the KSD and $P$-KSD through Lemma 2.
@@ A.4.1 Properties of $P$-KSD
Proof of Lemma 15  This follows directly from the symmetry of $P$-KSD (Shi and Mackey,, 2024, Lemma 1) and that the assumed conditions imply $\mathbb{E}_{\mathbf{X}\sim P}[(\mathcal{A}_{P}f)(\mathbf{X})]=0$ for all $f\in\mathcal{B}$ (Gorham and Mackey,, 2017, Proposition 1), so one of the two expectations in (48) vanishes when either of the two arguments coincides with $P$.
Barp et al., (2024, Theorem 1) shows that a KSD is equivalent to an MMD with the Stein kernel $u_{p}$. Since $P$-KSD is a KSD by Lemma 15, we can conclude that the $P$-KSD is also an MMD with kernel $u_{p}$, and that $P$-KSD has a double-expectation form similar to MMD.
@@
Proof of Lemma 16  The assumed conditions ensure $\mathbb{S}_{P}^{2}(Q,R)$ is well-defined. Since $P$-KSD is an MMD with reproducing kernel $u_{p}$, we can apply Gretton et al., (2012, Lemma 4) to conclude the first equality, and Gretton et al., (2012, Lemma 6) to yield the second.
@@
Let $R\in\mathcal{P}(\mathbb{R}^{d})$ and assume $\mathbb{E}_{\mathbf{Y}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y})]<\infty$. Further assume the conditions in Lemma 15 hold. For any $\theta>0$,
@@
Proof of Lemma 17  The $P$-KSD $\mathbb{S}_{P}(R,P)$ is well-defined under the assumed conditions, and hence so is $\mathbb{S}_{P}\big{(}\lambda R+(1-\lambda)P,P\big{)}$ for any $\lambda\in[0,1]$. Also since $P$-KSD is an MMD with (potentially unbounded) reproducing kernel $u_{p}$, the claim then follows by Sun and Zou, (2023, Proposition 3) and noting that their proof extends directly to potentially unbounded kernels.
@@
Assume $k$ is a tilted kernel satisfying the conditions in Lemma 2 and $P\in\mathcal{P}(\mathbb{R}^{d})$ has a density $p\in\mathcal{C}^{1}$. For any random sample $\mathbb{X}_{n}$ drawn from $Q\in\mathcal{P}(\mathbb{R}^{d})$ and any $\alpha>0$,
@@
Proof of Lemma 18  Let $Q_{n}$ denote the empirical distribution based on $\mathbb{X}_{n}$. By the first equality in Lemma 16, we can write $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\|\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X})]\|_{\mathcal{H}_{u}}$. Under the assumptions in Lemma 2, the function $u_{p}(\cdot,\mathbf{x}):\mathbb{R}^{d}\to\mathcal{H}_{u}$ is continuous for all $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $\|u_{p}(\cdot,\mathbf{x})\|_{\mathcal{H}_{u}}=u_{p}(\mathbf{x},\mathbf{x})$ by the reproducing property of the Stein kernel $u_{p}$ (see also the argument before (30)), and $\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})<\infty$ again by Lemma 2. We can hence apply the McDiarmid-type inequality for MMD in Tolstikhin et al., (2017, Proposition A.1) to conclude the claimed result.
@@ A.5 Validity of the Bootstrap Approach
We provide a proof for the validity of the bootstrap approach in Theorem 4. We first discuss the intuition, and then present two lemmas in Section A.5.1. The proof of Theorem 4 will be presented in Section A.5.2.
The bootstrap procedure used in robust-KSD is the same as the one in the standard KSD test, and it is not immediately obvious that this gives a valid decision threshold. For the standard KSD test, the V-statistic $D^{2}(\mathbb{X}_{n})$ is degenerate under the point null $H_{0}:Q=P$, i.e., $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,P)]=0$ when $Q=P$ (Liu et al.,, 2016, Theorem 4.1), and classic bootstrapping methods for degenerate V-statistics (Arcones and Gine,, 1992; Huskova and Janssen,, 1993) shows that the bootstrap quantile $q_{\infty,1-\alpha}$ is a valid decision threshold. However, the same argument cannot be applied directly to our robust test, because our null $H^{\mathrm{C}}_{0}$ is a composite one that contains not only $P$ but also other distributions $Q\neq P$, in which case $D^{2}(\mathbb{X}_{n})$ is no longer degenerate as shown by Liu et al., (2016, Theorem 4.1). Nevertheless, our proof shows that $q_{\infty,1-\alpha}$ is still a correct decision threshold.
@@
where $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$ is the $P$-KSD introduced in Section A.4. We will then prove that the RHS of the above inequality converges to the prescribed test level $\alpha$ as $n\to\infty$ by showing that $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is a valid bootstrap approximation for the $(1-\alpha)$-quantile of the distribution of $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$.
To see this, we first use the closed-form expression for $P$-KSD (Lemma 16) to write
@@
A naive alternative approach to bootstrap $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ is to note that it is degenerate for any $Q$ and use standard bootstrapping techniques for degenerate V-statistics (e.g., Arcones and Gine,, 1992, Theorem 3.5). However, this approach is not applicable here since, to compute the bootstrap sample, it requires the function $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ to be computable at any $\mathbf{x},\mathbf{x}^{\prime}$, which is not the case for $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ since it involves intractable expectations over $Q$ (recall that $u_{p}$ has mean zero under $P$, but not necessarily under $Q$).
@@ A.5.1 Bootstrapping the $P$-KSD Estimate
Our proof for Theorem 4 relies on Arcones and Gine, (1992, Lemma 3.4), which states that the asymptotic distribution of the second-order term in the Hoeffding’s decomposition of a V-statistic can be approximated by a bootstrap distribution. We restate this result for the case of $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ in Lemma 19.
@@
Let $\mathbb{X}_{\infty}=\{\mathbf{X}_{i}\}_{i=1}^{\infty}$ be a random sample where $\mathbf{X}_{i}\sim Q$ are independent, and for any $n$ define $\mathbb{X}_{n}\coloneqq\{\mathbf{X}_{i}\}_{i=1}^{n}$. Let $Q_{n}$ be the empirical measure based on $\mathbb{X}_{n}$ and define $\mathbb{X}_{n}^{\ast}=\{\mathbf{X}_{i}^{\ast}\}_{i=1}^{n}$, where $\mathbf{X}_{i}^{\ast}\sim Q_{n}$ are independent conditionally on $\mathbb{X}_{n}$. Assume $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$. Then, as $n\to\infty$,
@@
Assume $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$ and let $\mathbf{W}\coloneqq(W_{1},\ldots,W_{n})$ be a random vector distributed as $\textrm{Multinomial}(n;1/n,\ldots,1/n)$. Under the notation in Lemma 19, the following holds as $n\to\infty$,
@@
Proof of Lemma 20  As argued in (A.5), the squared $P$-KSD $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ can be written as a V-statistic with symmetric function $u_{p}(\cdot,\cdot;Q)$. Moreover, direct computation shows that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X};Q)]\equiv 0$, so the symmetric function $u_{p}(\cdot,\cdot;Q)$ is $Q$-degenerate of order 1 (Arcones and Gine,, 1992, pp. 5). On the other hand, it is well-known that the weighted bootstrap form can be equivalently written as an Efron’s resampled bootstrap statistic (Janssen,, 1994; Dehling and Mikosch,, 1994; Janssen,, 1997), i.e.,
@@
=$\stackrel{{\scriptstyle d}}{{=}}\;\frac{1}{n^{2}}\sum_{1\leq i,j\leq n}u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{X}_{j}^{\ast})-u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{X}_{j})-u_{p}(\mathbf{X}_{i},\mathbf{X}_{j}^{\ast})+u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$
$=\;\frac{1}{n^{2}}\sum_{1\leq i,j\leq n}u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{X}_{j}^{\ast};Q_{n})\;=\;\mathbb{S}_{P}^{2}(\mathbb{X}_{n}^{\ast},\mathbb{X}_{n})\;,$
where $\mathbf{X}_{i}^{\ast}$ and $Q_{n}$ are defined in Lemma 19, the notation =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution, the second last line follows by the definition of $u_{p}(\cdot,\cdot;Q_{n})$, and the last line. Under the assumed moment condition, we can apply Lemma 19 together with the above derivation to conclude that a version of (50) with the squared statistics holds, i.e.,
@@
The claim then follows by noting that the mapping $u\mapsto\sqrt{u}$ is everywhere continuous on $[0,\infty)$ and that weak convergence is preserved by continuous function by the Continuous Mapping Theorem (Van der Vaart,, 2000, Theorem 2.3).
A.5.2 Proof of Theorem 4
We first show that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$, so in particular $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})]<\infty$ and we can apply Lemma 17. Defining $\mathbf{s}_{p,w}(x)=w(x)\mathbf{s}_{p}(x)$, we have

### v4

@@ A.4 $P$-Targeted Kernel Stein Discrepancy
This section states preliminary results which will be needed to prove Theorem˜4. The main tool we use is the following generalized definition of KSD, which was originally proposed in Shi and Mackey, (2024, Definition 4).
@@
  • Lemma˜15 shows that the standard KSD can be written as a $P$-KSD.
  • Lemma˜16 shows that $\mathbb{S}_{P}(Q,R)$ is equivalent to an MMD, and that it admits a closed form involving expectations over $Q$ and $R$.
  • Lemma˜17 shows that the $P$-KSD-projection of a probability measure onto a (standard) KSD-ball centered at $P$ admits a closed-form expression.
  • Lemma˜18 restates Tolstikhin et al., (2017, Proposition A.1), which is a McDiarmid-type inequality for MMD. It also applies to $P$-KSD, as an $P$-KSD is also an MMD. We will use it in Appendix˜D to derive a robust-KSD test that is well-calibrated for finite samples. Other deviation bounds instead of the McDiarmid bound could also be used to construct a similar test;see Remark˜24.1 for a discussion.
The MMD has already been studied extensively in the context of both Bayesian and frequentist robust parameter estimation; see, e.g., Briol et al., (2019); Chérief-Abdellatif and Alquier, (2020, 2022); Dellaporta et al., (2022); Dellaporta and Damoulas, (2023); Alquier and Gerber, (2024). A key assumption to ensure robustness in all of these papers is that the kernel is assumed to be bounded, and this is exactly what we are able to achieve with the KSD and $P$-KSD through Lemma˜2.
@@ A.4.1 Properties of $P$-KSD
Proof of Lemma˜15  This follows directly from the symmetry of $P$-KSD (Shi and Mackey,, 2024, Lemma 1) and that the assumed conditions imply $\mathbb{E}_{\mathbf{X}\sim P}[(\mathcal{A}_{P}f)(\mathbf{X})]=0$ for all $f\in\mathcal{B}$ (Gorham and Mackey,, 2017, Proposition 1), so one of the two expectations in (48) vanishes when either of the two arguments coincides with $P$.
Barp et al., (2024, Theorem 1) shows that a KSD is equivalent to an MMD with the Stein kernel $u_{p}$. Since $P$-KSD is a KSD by Lemma˜15, we can conclude that the $P$-KSD is also an MMD with kernel $u_{p}$, and that $P$-KSD has a double-expectation form similar to MMD.
@@
Proof of Lemma˜16  The assumed conditions ensure $\mathbb{S}_{P}^{2}(Q,R)$ is well-defined. Since $P$-KSD is an MMD with reproducing kernel $u_{p}$, we can apply Gretton et al., (2012, Lemma 4) to conclude the first equality, and Gretton et al., (2012, Lemma 6) to yield the second.
@@
Let $R\in\mathcal{P}(\mathbb{R}^{d})$ and assume $\mathbb{E}_{\mathbf{Y}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y})]<\infty$. Further assume the conditions in Lemma˜15 hold. For any $\theta>0$,
@@
Proof of Lemma˜17  The $P$-KSD $\mathbb{S}_{P}(R,P)$ is well-defined under the assumed conditions, and hence so is $\mathbb{S}_{P}\big{(}\lambda R+(1-\lambda)P,P\big{)}$ for any $\lambda\in[0,1]$. Also since $P$-KSD is an MMD with (potentially unbounded) reproducing kernel $u_{p}$, the claim then follows by Sun and Zou, (2023, Proposition 3) and noting that their proof extends directly to potentially unbounded kernels.
@@
Assume $k$ is a tilted kernel satisfying the conditions in Lemma˜2 and $P\in\mathcal{P}(\mathbb{R}^{d})$ has a density $p\in\mathcal{C}^{1}$. For any random sample $\mathbb{X}_{n}$ drawn from $Q\in\mathcal{P}(\mathbb{R}^{d})$ and any $\alpha>0$,
@@
Proof of Lemma˜18  Let $Q_{n}$ denote the empirical distribution based on $\mathbb{X}_{n}$. By the first equality in Lemma˜16, we can write $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\|\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X})]\|_{\mathcal{H}_{u}}$. Under the assumptions in Lemma˜2, the function $u_{p}(\cdot,\mathbf{x}):\mathbb{R}^{d}\to\mathcal{H}_{u}$ is continuous for all $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $\|u_{p}(\cdot,\mathbf{x})\|_{\mathcal{H}_{u}}=u_{p}(\mathbf{x},\mathbf{x})$ by the reproducing property of the Stein kernel $u_{p}$ (see also the argument before (30)), and $\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})<\infty$ again by Lemma˜2. We can hence apply the McDiarmid-type inequality for MMD in Tolstikhin et al., (2017, Proposition A.1) to conclude the claimed result.
@@ A.5 Validity of the Bootstrap Approach
We provide a proof for the validity of the bootstrap approach in Theorem˜4. We first discuss the intuition, and then present two lemmas in Section˜A.5.1. The proof of Theorem˜4 will be presented in Section˜A.5.2.
The bootstrap procedure used in robust-KSD is the same as the one in the standard KSD test, and it is not immediately obvious that this gives a valid decision threshold. For the standard KSD test, the V-statistic $D^{2}(\mathbb{X}_{n})$ is degenerate under the point null $H_{0}:Q=P$, i.e., $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,P)]=0$ when $Q=P$ (Liu et al.,, 2016, Theorem 4.1), and classic bootstrapping methods for degenerate V-statistics (Arcones and Gine,, 1992; Huskova and Janssen,, 1993) show that the bootstrap quantile $q_{\infty,1-\alpha}$ is a valid decision threshold. However, the same argument cannot be applied directly to our robust test, because our null $H^{\mathrm{C}}_{0}$ is a composite one that contains not only $P$ but also other distributions $Q\neq P$, in which case $D^{2}(\mathbb{X}_{n})$ is no longer degenerate as shown by Liu et al., (2016, Theorem 4.1). Nevertheless, our proof shows that $q_{\infty,1-\alpha}$ is still a correct decision threshold.
@@
where $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$ is the $P$-KSD introduced in Section˜A.4. We will then prove that the RHS of the above inequality converges to the prescribed test level $\alpha$ as $n\to\infty$ by showing that $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is a valid bootstrap approximation for the $(1-\alpha)$-quantile of the distribution of $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$.
To see this, we first use the closed-form expression for $P$-KSD (Lemma˜16) to write
@@
A naive alternative approach to bootstrap $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ is to note that it is degenerate for any $Q$ and use standard bootstrapping techniques for degenerate V-statistics (e.g., Arcones and Gine,, 1992, Theorem 3.5). However, this approach is not applicable here since, to compute the bootstrap sample, it requires the function $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ to be computable at any $\mathbf{x},\mathbf{x}^{\prime}$, which is not the case since $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ involves intractable expectations over $Q$ (recall that $u_{p}$ has mean zero under $P$, but not necessarily under $Q$).
@@ A.5.1 Bootstrapping the $P$-KSD Estimate
Our proof for Theorem˜4 relies on Arcones and Gine, (1992, Lemma 3.4), which states that the asymptotic distribution of the second-order term in the Hoeffding’s decomposition of a V-statistic can be approximated by a bootstrap distribution. We restate this result for the case of $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ in Lemma˜19.
@@
Let $\mathbb{X}_{\infty}=\{\mathbf{X}_{i}\}_{i=1}^{\infty}$ be a random sample where $\mathbf{X}_{i}\sim Q$ are independent, and for any $n$ define $\mathbb{X}_{n}\coloneqq\{\mathbf{X}_{i}\}_{i=1}^{n}$. Let $Q_{n}$ be the empirical measure based on $\mathbb{X}_{n}$ and define $\mathbb{X}_{n}^{\ast}=\{\mathbf{X}_{i}^{\ast}\}_{i=1}^{n}$, where $\mathbf{X}_{i}^{\ast}\sim Q_{n}$ are independent conditionally on $\mathbb{X}_{n}$. Assume $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$. Then, as $n\to\infty$, the following holds almost surely,
@@
Assume $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$ and let $\mathbf{W}\coloneqq(W_{1},\ldots,W_{n})$ be a random vector distributed as $\textrm{Multinomial}(n;1/n,\ldots,1/n)$. Under the notation in Lemma˜19, the following holds as $n\to\infty$,
@@
Proof of Lemma˜20  As argued in (A.5), the squared $P$-KSD $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ can be written as a V-statistic with symmetric function $u_{p}(\cdot,\cdot;Q)$. Moreover, direct computation shows that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X};Q)]\equiv 0$, so the symmetric function $u_{p}(\cdot,\cdot;Q)$ is $Q$-degenerate of order 1 (Arcones and Gine,, 1992, pp. 5). On the other hand, it is well-known that the weighted bootstrap form can be equivalently written as an Efron’s resampled bootstrap statistic (Janssen,, 1994; Dehling and Mikosch,, 1994; Janssen,, 1997), i.e.,
@@
$\stackrel{{\scriptstyle d}}{{=}}\;\frac{1}{n^{2}}\sum_{1\leq i,j\leq n}u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{X}_{j}^{\ast})-u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{X}_{j})-u_{p}(\mathbf{X}_{i},\mathbf{X}_{j}^{\ast})+u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$
$=\;\frac{1}{n^{2}}\sum_{1\leq i,j\leq n}u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{X}_{j}^{\ast};Q_{n})$
$=\;\mathbb{S}_{P}^{2}(\mathbb{X}_{n}^{\ast},\mathbb{X}_{n})\;,$
where $\mathbf{X}_{i}^{\ast}$ and $Q_{n}$ are defined in Lemma˜19, the notation $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution, the second last line follows by the definition of $u_{p}(\cdot,\cdot;Q_{n})$, and the last line follows from Lemma˜16. Under the assumed moment condition, we can apply Lemma˜19 together with the above derivation to conclude that a version of (50) with the squared statistics holds, i.e.,
@@
The claim then follows by noting that the mapping $u\mapsto\sqrt{u}$ is everywhere continuous on $0,\infty)$ and that weak convergence is preserved by continuous function by the Continuous Mapping Theorem (Van der Vaart,, 2000, Theorem 2.3).
A.5.2 Proof of Theorem˜4
Proof of Theorem˜4  We first show that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$, so in particular $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})]<\infty$ and we can apply Lemma˜17. Defining $\mathbf{s}_{p,w}(x)=w(x)\mathbf{s}_{p}(x)$, we have